Find the number of particles each carrying the charge of an electron that passes between the terminals of a 12-V car battery when a 60-W headlight is on for one hour.




With Complete Solution. Thank you.

1
Expert's answer
2022-02-08T08:16:22-0500

Given:

"V=12\:\rm V"

"P=60\:\rm W"

"t=3600\:\rm s"


The work done by electric field is given by

"W=qV=Pt"

The charge that passes between the terminals of a battery

"q=\frac{Pt}{V}=\frac{60*3600}{12}=18000 \: \rm C"

Hence, the number of particles

"n=\frac{q}{e}=\frac{1.8*10^4}{1.6*10^{-19}}=1.125*10^{23}"